OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4.

OpenStudy (anonymous):

let m be the slope pf the required line. we know that if two lines are perpendicular, then the product of their slopes is -1. so -1/2m=-1. therefore m=2. now the equation of the line passing through (-2,4) and having slope 2 is given by y-4=2(x+2) which when simplified gives y-2x=8

OpenStudy (anonymous):

y = -2x - 8

OpenStudy (anonymous):

is that the answer